# On von Neumann regularity of ample groupoid algebras

### v2

@@ 4.1. Graded regularity
Let $K$ be a field, $R$ a $K$-algebra and $H$ a group with identity $e$. Then $R$ is $H$-graded if $R$ has a direct sum decomposition $R=\bigoplus_{h\in H}R_{h}$, where each $R_{h}$ is a $K$-vector subspace of $R$ such that $R_{h}R_{k}\subseteq R_{hk}$. The summand $R_{h}$ is called the homogeneous component of degree $h$. The homogeneous component $R_{e}$ is a $K$-subalgebra of $R$. An element is homogenous if it belongs to a homogeneous component.
@@
A graded ring $R$ is graded (von Neumann) regular if for every homogeneous element $x\in R$ there exists $y\in R$ such that $x=xyx$.
@@
Let $\mathscr{G}$ be an ample groupoid and let $H$ be a discrete group. A cocycle from $\mathscr{G}$ to $H$ is a map $c\colon\mathscr{G}\to H$ such that $c(\gamma^{-1})=c(\gamma)^{-1}$, for all $\gamma\in\mathscr{G}$, and $c(\gamma\eta)=c(\gamma)c(\eta)$ for all $(\gamma,\eta)\in\mathscr{G}^{(2)}$. A continuous cocycle $c\colon\mathscr{G}\to H$ induces an an $H$-grading $\{(K\mathscr{G})_{h}\}_{h}$ of $K\mathscr{G}$. This is done in detail in [9] for Hausdorff groupoids. The details for non-Hausdorff groupoids are similar, but we present them for completeness. Call a compact open bisection $U\in\Gamma_{c}(\mathscr{G})$ homogenous of degree $h\in H$ if $U\subseteq c^{-1}(h)$ for some $h\in H$. Then $(K\mathscr{G})_{h}$ is the $K$-span of the characteristic functions of homogeneous compact open bisections of degree $h$. Note that each $c^{-1}(h)$ with $h\in H$ is clopen, and so if $U$ is any compact open bisection, then $U=\bigcup_{h\in H}(U\cap c^{-1}(h))$ and the $U\cap c^{-1}(h)$ are pairwise disjoint homogeneuous compact open bisections. By compactness of $U$, only finitely many of these intersections, say corresponding to $h_{1},\ldots,h_{n}$, are nonempty, and so $1_{U}=\sum_{i=1}^{n}1_{(c^{-1}(h_{i})\cap U)}$, showing that $\sum_{h\in H}(K\mathscr{G})_{h}=K\mathscr{G}$. The sum is direct since the $c^{-1}(h)$ are pairwise disjoint and elements of $(K\mathscr{G})_{h}$ are supported on $c^{-1}(h)$. One can, in fact, show that $(K\mathscr{G})_{h}$ consists of those elements of $K\mathscr{G}$ supported on $c^{-1}(h)$, but we shall never need this description. This argument also shows that the homogeneous compact open bisections form a basis for the topology of $\mathscr{G}$.
@@
Notice that $\mathscr{G}^{(0)}\subseteq c^{-1}(e)=\mathscr{G}_{e}$ and that $\mathscr{G}_{e}$ is a clopen subgroupoid and hence is ample. Moreover, $(K\mathscr{G})_{e}=K\mathscr{G}_{e}$ by definition. From now on, we write $K\mathscr{G}_{h}$ instead of the more cumbersome $(K\mathscr{G})_{h}$.
@@
The following is [21, Definition 2.9]. Let $R=\bigoplus_{h\in H}R_{h}$ be an $H$-graded $K$-algebra. If, for every $h\in H$ and $r\in R_{h}$, there exist $\epsilon_{h}(r)\in R_{h}R_{h^{-1}}$ and $\epsilon_{h}^{\prime}(r)\in R_{h^{-1}}R_{h}$ such that $\epsilon_{g}(r)r=r=r\epsilon_{g}^{\prime}(r)$, then $R$ is called nearly epsilon-strongly $H$-graded. It is shown in [21] that being nearly epsilon-strongly $H$-graded is a necessary condition for graded regularity.
@@
Let $\mathscr{G}$ be an ample groupoid with a continuous cocycle $c\colon\mathscr{G}\to H$ and let $K$ be a field. Then $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded.
@@
Let $\mathscr{G}=\bigoplus_{h\in H}K\mathscr{G}_{h}$ be the $H$-grading induced by the continuous cocycle $c\colon\mathscr{G}\to H$ as above.
@@
$\epsilon_{h}(1_{U})=1_{UU^{-1}}=1_{U}1_{U^{-1}}\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}},$
@@
Similarly, defining $\epsilon^{\prime}_{h}(1_{U})=1_{U^{-1}U}$ yields that $\epsilon^{\prime}_{h}(1_{U})\in K\mathscr{G}_{h^{-1}}\,K\mathscr{G}_{h}$ and $1_{U}\epsilon^{\prime}_{h}(1_{U})=1_{U}$.
Now consider an arbitrary $f\in K\mathscr{G}_{h}$. Then there are compact open bisections $U_{1},U_{2},\ldots,U_{n}$, with each $U_{i}\subseteq c^{-1}(h)$, and $k_{1},k_{2},\cdots,k_{n}\in K$ such that $f=\sum_{i=1}^{n}k_{i}1_{U_{i}}$. In particular, $\epsilon_{h}(1_{U_{i}})\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}}$ for each $i$. Put $U=\bigcup_{i=1}^{n}U_{i}U_{i}^{-1}$. Then $U$ is a compact open set in $\mathscr{G}^{(0)}$. Put $\epsilon_{h}(f)=1_{U}$. Since $U\subseteq\mathscr{G}^{(0)}$ and $\mathscr{G}^{(0)}$ is Hausdorff, it follows from the inclusion-exclusion principle that $1_{U}=\epsilon_{h}(f)$ is an integer combination of the $1_{U_{i}U_{i}^{-1}}$, and hence $\epsilon_{h}(f)\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}}$. Similarly, putting $U^{\prime}=\bigcup_{i=1}^{n}U_{i}^{-1}U_{i}$ and defining $\epsilon^{\prime}_{h}(f)=1_{U^{\prime}}$, we have that $\epsilon^{\prime}_{h}(f)\in K\mathscr{G}_{h^{-1}}\,K\mathscr{G}_{h}$. It is straightforward from the construction that $\epsilon_{h}(f)f=f=f\epsilon^{\prime}_{h}(f)$. Hence, $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded. ∎
@@
Let $\mathscr{G}$ be an ample groupoid, let $K$ be a field, let $H$ be a discrete group and let $c\colon\mathscr{G}\to H$ be a continuous cocycle. Then $K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}$ is regular.
@@
By Proposition 4.1, $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded. Thus, by [21, Theorem 1.2], $K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}$ is regular. ∎
Theorem 4.2 allows us to upgrade Theorems 3.7 and 3.12 to a theorem about graded regularity.
@@
Let $\mathscr{G}$ be an ample groupoid, let $K$ be a field, let $H$ be a discrete group and let $c\colon\mathscr{G}\to H$ be a continuous cocycle.
@@
    If $\mathscr{G}_{e}$ is approximately quasi-compact and the characteristic of $K$ divides the order of no isotropy element of $\mathscr{G}_{e}$, then $K\mathscr{G}$ is graded regular with respect to the $H$-grading induced by $c$.
@@
    If $K$ is of characteristic 00, then $K\mathscr{G}$ is graded regular (with respect to the $H$-grading induced by $c$) if and only if $\mathscr{G}_{e}$ is approximately quasi-compact.
@@ 5. Applications
Our first application is to recover the nontrivial part of Okniński’s result on regularity of inverse semigroups in characteristic zero from Theorem 3.7.
Theorem 5.1 (Okniński).
Let $S$ be an inverse semigroup and $K$ a field of characteristic zero. Then $KS$ is regular if and only if $S$ is locally finite.
Proof.
We already proved in Proposition 2.5 that if $S$ is locally finite, then $KS$ is regular. For the converse, assume that $KS$ is regular. Let $\mathscr{G}_{S}$ denote the universal groupoid of $S$. Since $KS\cong K\mathscr{G}_{S}$, we have that $K\mathscr{G}_{S}$ is regular. Then, by Theorem 3.7, we have that $\mathscr{G}_{S}$ is approximately quasi-compact, and hence $S$ is locally finite by Corollary 3.5 since the canonical mapping $\psi\colon S\to\Gamma_{c}(\mathscr{G}_{S})$ given by $\psi(s)=(s,D_{s^{*}s})$ is injective. ∎
Next we give an example to show that the necessary conditions in Theorem 3.1 are not sufficient to guarantee regularity of the algebra of a groupoid by giving examples of principal groupoids with nonregular algebras. Recall that if a discrete group $G$ acts on a locally compact, Hausdorff and totally disconnected space $X$ by homeomorphisms, then the groupoid of germs of this action is just the so-called transformation groupoid $\mathscr{G}$ with underlying space $G\times X$ (with the product topology) and unit space $\{1\}\times X$ (identified with $X$) with $\mathop{\boldsymbol{s}}(g,x)=x$ and $\mathop{\boldsymbol{r}}(g,x)=gx$. Products are given by $(h,gx)(g,x)=(hg,x)$ and $(g,x)^{-1}=(g^{-1},gx)$. The isotropy group of the transformation groupoid at $x\in X$ is just stabilizer of $x$ in $G$. In particular, if the action of $G$ is free, then the transformation groupoid is principal. It is well known that $K\mathscr{G}$ is isomorphic to the skew group ring $C_{c}(X,K)\rtimes G$ where $C_{c}(X,K)$ is the ring of compactly supported locally constant functions $X\to K$ and $G$ acts on this ring via $(gf)(x)=f(g^{-1}x)$. The first part of the following result extends a result in [18].
@@
    If $K$ is a field of characteristic 00, then $K\mathscr{G}$ is regular if and only if $G$ is locally finite.
@@
The first item follows from Corollary 3.5 since the map $\psi\colon G\to\Gamma_{c}(\mathscr{G})$ given by $\psi(g)=\{g\}\times X$ is injective. The second item follows from the first and Theorem 3.7. ∎
For example, if we consider a minimal (nonperiodic) action of $\mathbb{Z}$ on the Cantor set $X$ (e.g., a nonperiodic minimal shift space), then the associated transformation groupoid $\mathscr{G}$ is a principal groupoid (and hence trivially satisfies the necessary conditions in Theorem 3.1), but $\mathbb{C}\mathscr{G}$ is not regular.
@@ 5.1. Leavitt path algebras
Then the collection $\{Z(\mu,F,\nu)\}$, ranging over $\mu,\nu\in E^{*}$ with $r(\mu)=r(\nu)$ and finite $F\subseteq\{e\in E^{1}\mid s(e)=r(\mu)\}$, forms a basis of compact open bisections for a locally compact Hausdorff topology on $\mathscr{G}_{E}$ [26, Lemma 9.2]. In particular, $\mathscr{G}_{E}$ is an ample groupoid. The unit space $\mathscr{G}_{E}^{(0)}$ is identified with $\partial E$.
@@
Let $K$ be a field and let $L_{K}(E)$ denote the Leavitt path algebra of $E$ over $K$ (see [2] for the definition of $L_{K}(E)$). Then $L_{K}(E)\cong K\mathscr{G}_{E}$ with the isomorphism given by $\mu\nu^{*}\mapsto 1_{Z(\mu,\nu)}$ [9, Proposition 4.3]. In [3, Theorem 1] it is proved that $L_{K}(E)$ is regular if and only if $E$ acyclic. Using our groupoid characterization, we can give a short proof of this equivalence, and expand on it.
@@
Let $E$ be a directed graph and $K$ a field. The following are equivalent.
@@
We already observed that acyclic is equivalent to principal. If $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular, then $\mathscr{G}_{E}$ has locally finite isotropy by Theorem 3.1 and hence $E$ must be principal as $\mathscr{G}_{E}$ has no infinite cyclic isotropy group. Suppose now that $E$ is acyclic. We show that $\mathscr{G}_{E}$ is approximately quasi-compact.
@@
If $\mathscr{G}_{E}$ is approximately quasi-compact (and hence approximately uniformly bounded), then all its isotropy groups must be locally finite. Since the isotropy groups of $\mathscr{G}_{E}$ are always subgroups of $\mathbb{Z}$, they must then be trivial and so $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular by Theorem 3.12. ∎
Next we show that we can recover [20, Theorem 10], which states that the Leavitt path algebra of an arbitrary graph is graded von Neumann regular. Every Leavitt path algebra $L_{K}(E)$ over a field $K$ is $\mathbb{Z}$-graded, with the grading given by homogeneous components
@@
Let $E$ be an arbitrary graph and $K$ a field. Then the Leavitt path algebra $L_{K}(E)$ is a graded von Neumann regular ring.
@@
$\mathscr{G}_{0}(F,n)=\bigcup_{\alpha,\beta\in F^{*},\mathop{\boldsymbol{r}}(\alpha)=\mathop{\boldsymbol{r}}(\beta),|\alpha|=|\beta|\leq n}Z(\alpha,\emptyset,\beta).$

### v3

@@ 4.1. Graded regularity
If $H$ is a group with identity $e$, then a ring $\Lambda$ is $H$-graded if $\Lambda$ has a direct sum decomposition $\Lambda=\bigoplus_{h\in H}\Lambda_{h}$, where each $\Lambda_{h}$ is an additive subgroup of $\Lambda$ such that $\Lambda_{h}\Lambda_{k}\subseteq\Lambda_{hk}$ for all $h,k\in H$. The summand $\Lambda_{h}$ is called the homogeneous component of degree $h$. The homogeneous component $\Lambda_{e}$ is a subring of $\Lambda$. An element is homogenous if it belongs to a homogeneous component.
@@
A graded ring $\Lambda$ is graded (von Neumann) regular if for every homogeneous element $x\in\Lambda$ there exists $y\in\Lambda$ such that $x=xyx$.
@@
Let $\mathscr{G}$ be an ample groupoid and let $H$ be a discrete group. A cocycle from $\mathscr{G}$ to $H$ is a map $c\colon\mathscr{G}\to H$ such that $c(\gamma\eta)=c(\gamma)c(\eta)$ for all $(\gamma,\eta)\in\mathscr{G}^{(2)}$. A continuous cocycle $c\colon\mathscr{G}\to H$ induces an an $H$-grading $\{(R\mathscr{G})_{h}\}_{h}$ of $R\mathscr{G}$. This is done in detail in [9] for Hausdorff groupoids. The details for non-Hausdorff groupoids are similar, but we present them for completeness. Call a compact open bisection $U\in\Gamma_{c}(\mathscr{G})$ homogenous of degree $h\in H$ if $U\subseteq c^{-1}(h)$ for some $h\in H$. Then $(R\mathscr{G})_{h}$ is the $R$-span of the characteristic functions of homogeneous compact open bisections of degree $h$. Note that each $c^{-1}(h)$ with $h\in H$ is clopen, and so if $U$ is any compact open bisection, then $U=\bigcup_{h\in H}(U\cap c^{-1}(h))$ and the $U\cap c^{-1}(h)$ are pairwise disjoint homogeneous compact open bisections. By compactness of $U$, only finitely many of these intersections, say corresponding to $h_{1},\ldots,h_{n}$, are nonempty, and so $1_{U}=\sum_{i=1}^{n}1_{(c^{-1}(h_{i})\cap U)}$, showing that $\sum_{h\in H}(R\mathscr{G})_{h}=R\mathscr{G}$. The sum is direct since the $c^{-1}(h)$ are pairwise disjoint and elements of $(R\mathscr{G})_{h}$ are supported on $c^{-1}(h)$. One can, in fact, show that $(R\mathscr{G})_{h}$ consists of those elements of $R\mathscr{G}$ supported on $c^{-1}(h)$, but we shall never need this description. This argument also shows that the homogeneous compact open bisections form a basis for the topology of $\mathscr{G}$.
@@
Notice that $\mathscr{G}^{(0)}\subseteq c^{-1}(e)=\mathscr{G}_{e}$ and that $\mathscr{G}_{e}$ is a clopen subgroupoid and hence is ample. Moreover, $(R\mathscr{G})_{e}=R\mathscr{G}_{e}$ by definition. From now on, we write $R\mathscr{G}_{h}$ instead of the more cumbersome $(R\mathscr{G})_{h}$.
@@
The following is [22, Definition 2.9]. Let $\Lambda=\bigoplus_{h\in H}\Lambda_{h}$ be an $H$-graded $K$-algebra. If, for every $h\in H$ and $r\in\Lambda_{h}$, there exist $\epsilon_{h}(r)\in\Lambda_{h}\Lambda_{h^{-1}}$ and $\epsilon_{h}^{\prime}(r)\in\Lambda_{h^{-1}}\Lambda_{h}$ such that $\epsilon_{g}(r)r=r=r\epsilon_{g}^{\prime}(r)$, then $\Lambda$ is called nearly epsilon-strongly $H$-graded. It is shown in [22] that being nearly epsilon-strongly $H$-graded is a necessary condition for graded regularity.
@@
Let $\mathscr{G}$ be an ample groupoid with a continuous cocycle $c\colon\mathscr{G}\to H$ and let $R$ be a ring. Then $R\mathscr{G}$ is nearly epsilon-strongly $H$-graded.
@@
Let $\mathscr{G}=\bigoplus_{h\in H}R\mathscr{G}_{h}$ be the $H$-grading induced by the continuous cocycle $c\colon\mathscr{G}\to H$ as above.
@@
$\epsilon_{h}(1_{U})=1_{UU^{-1}}=1_{U}1_{U^{-1}}\in R\mathscr{G}_{h}\,R\mathscr{G}_{h^{-1}},$
@@
Similarly, defining $\epsilon^{\prime}_{h}(1_{U})=1_{U^{-1}U}$ yields that $\epsilon^{\prime}_{h}(1_{U})\in R\mathscr{G}_{h^{-1}}\,R\mathscr{G}_{h}$ and $1_{U}\epsilon^{\prime}_{h}(1_{U})=1_{U}$.
Now consider an arbitrary $f\in R\mathscr{G}_{h}$. Then there are compact open bisections $U_{1},U_{2},\ldots,U_{n}$, with each $U_{i}\subseteq c^{-1}(h)$, and $r_{1},r_{2},\cdots,r_{n}\in R$ such that $f=\sum_{i=1}^{n}r_{i}1_{U_{i}}$. In particular, $\epsilon_{h}(1_{U_{i}})\in R\mathscr{G}_{h}\,R\mathscr{G}_{h^{-1}}$ for each $i$. Put $U=\bigcup_{i=1}^{n}U_{i}U_{i}^{-1}$. Then $U$ is a compact open set in $\mathscr{G}^{(0)}$. Put $\epsilon_{h}(f)=1_{U}$. Since $U\subseteq\mathscr{G}^{(0)}$ and $\mathscr{G}^{(0)}$ is Hausdorff, it follows from the inclusion-exclusion principle that $1_{U}=\epsilon_{h}(f)$ is an integer combination of the $1_{U_{i}U_{i}^{-1}}$, and hence $\epsilon_{h}(f)\in R\mathscr{G}_{h}\,R\mathscr{G}_{h^{-1}}$. Similarly, putting $U^{\prime}=\bigcup_{i=1}^{n}U_{i}^{-1}U_{i}$ and defining $\epsilon^{\prime}_{h}(f)=1_{U^{\prime}}$, we have that $\epsilon^{\prime}_{h}(f)\in R\mathscr{G}_{h^{-1}}\,R\mathscr{G}_{h}$. It is straightforward from the construction that $\epsilon_{h}(f)f=f=f\epsilon^{\prime}_{h}(f)$. Hence, $R\mathscr{G}$ is nearly epsilon-strongly $H$-graded. ∎
@@
Let $\mathscr{G}$ be an ample groupoid, $R$ a unital ring, $H$ a discrete group and $c\colon\mathscr{G}\to H$ a continuous cocycle. Then $R\mathscr{G}$ is graded regular if and only if $R\mathscr{G}_{e}$ is regular.
@@
By Proposition 4.1, $R\mathscr{G}$ is nearly epsilon-strongly $H$-graded. Thus, by [22, Theorem 1.2], $R\mathscr{G}$ is graded regular if and only if $R\mathscr{G}_{e}$ is regular. ∎
Theorem 4.2 allows us to upgrade Theorem A to a theorem about graded regularity.
@@
Let $\mathscr{G}$ be an ample groupoid, $R$ a unital ring, $H$ a discrete group and $c\colon\mathscr{G}\to H$ a continuous cocycle. Then $R\mathscr{G}$ is graded von Neumann regular (with respect to the $H$-grading induced by $c$) if and only if the following conditions hold:
@@
    $R$ is regular
@@
    $\mathscr{G}_{e}$ is approximately quasi-compact
  3. (3)
    the order of each finite subgroup of an isotropy group of $\mathscr{G}_{e}$ is invertible in $R$.
@@ 5. Applications
Our first application is to prove Theorem B.
Proof of Theorem B.
The sufficiency of the conditions is proved in Proposition 2.4. For sufficiency, suppose that $RS$ is regular. Let $\mathscr{G}_{S}$ denote the universal groupoid of $S$. Then $R\mathscr{G}_{S}\cong RS$ is regular, and so $R$ is regular, $\mathscr{G}_{S}$ is approximately quasi-compact and the order of each finite subgroup of an isotropy group of $\mathscr{G}_{S}$ is invertible in $R$. Then $S$ is locally finite by Corollary 3.5 since the canonical mapping $\psi\colon S\to\Gamma_{c}(\mathscr{G}_{S})$ given by $\psi(s)=(s,D_{s^{*}s})$ is injective. A maximal subgroup $G_{e}$ of $S$ is isomorphic to the isotropy group at the principal character $x_{e}$, and hence the order of any finite subgroup of $G_{e}$ is invertible in $R$. ∎
Next we give an example to show that the necessary conditions for regularity in [4] are not sufficient to guarantee regularity of the algebra of a groupoid by giving examples of principal groupoids with nonregular algebras over any regular unital ring $R$. Recall that if a discrete group $G$ acts on a locally compact, Hausdorff and totally disconnected space $X$ by homeomorphisms, then the groupoid of germs of this action is just the so-called transformation groupoid $\mathscr{G}$ with underlying space $G\times X$ (with the product topology) and unit space $\{1\}\times X$ (identified with $X$) with $\mathop{\boldsymbol{s}}(g,x)=x$ and $\mathop{\boldsymbol{r}}\nolimits(g,x)=gx$. Products are given by $(h,gx)(g,x)=(hg,x)$ and $(g,x)^{-1}=(g^{-1},gx)$. The isotropy group of the transformation groupoid at $x\in X$ is just stabilizer of $x$ in $G$. In particular, if the action of $G$ is free, then the transformation groupoid is principal. It is well known that $K\mathscr{G}$ is isomorphic to the skew group ring $C_{c}(X,R)\rtimes G$ where $C_{c}(X,R)$ is the ring of compactly supported locally constant functions $X\to R$ and $G$ acts on this ring via $(gf)(x)=f(g^{-1}x)$. The first part of the following result extends a result in [18].
@@
    If $R$ is a regular unital ring, then $R\mathscr{G}$ is regular if and only if $G$ is locally finite and the order of each finite subgroup of an isotropy group of the action is invertible in $R$.
@@
The first item follows from Corollary 3.5 since the map $\psi\colon G\to\Gamma_{c}(\mathscr{G})$ given by $\psi(g)=\{g\}\times X$ is injective. The second item follows from the first and Theorem A. ∎
For example, if we consider a minimal (nonperiodic) action of $\mathbb{Z}$ on the Cantor set $X$ (e.g., a nonperiodic minimal shift space), then the associated transformation groupoid $\mathscr{G}$ is a principal groupoid (and hence trivially satisfies the necessary conditions in [4]), but $\mathbb{C}\mathscr{G}$ is not regular.
@@ 5.1. Leavitt path algebras
Then the collection $\{Z(\mu,F,\nu)\}$, ranging over $\mu,\nu\in E^{*}$ with $r(\mu)=r(\nu)$ and finite $F\subseteq\{e\in E^{1}\mid s(e)=r(\mu)\}$, forms a basis of compact open bisections for a locally compact Hausdorff topology on $\mathscr{G}_{E}$ [28, Lemma 9.2]. In particular, $\mathscr{G}_{E}$ is an ample groupoid. The unit space $\mathscr{G}_{E}^{(0)}$ is identified with $\partial E$.
@@
Let $K$ be a commutative ring with unit and let $L_{K}(E)$ denote the Leavitt path algebra of $E$ over $K$ (see [2] for the definition of $L_{K}(E)$). Then $L_{K}(E)\cong K\mathscr{G}_{E}$ with the isomorphism given by $\mu\nu^{*}\mapsto 1_{Z(\mu,\nu)}$ [9, Proposition 4.3]. In [3, Theorem 1] it is proved that $L_{K}(E)$ is regular if and only if $E$ acyclic. Using our groupoid characterization, we can give a short proof of this equivalence, and expand on it.
@@
Let $E$ be a directed graph and $K$ a von Neumann regular commutative ring. The following are equivalent.
@@
We already observed that acyclic is equivalent to principal. If $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular, then $\mathscr{G}_{E}$ has locally finite isotropy by Theorem A and hence $E$ must be principal, as $\mathscr{G}_{E}$ has no infinite cyclic isotropy group. Suppose now that $E$ is acyclic. We show that $\mathscr{G}_{E}$ is approximately quasi-compact.
@@
If $\mathscr{G}_{E}$ is approximately quasi-compact (and hence approximately uniformly bounded), then all its isotropy groups must be locally finite. Since the isotropy groups of $\mathscr{G}_{E}$ are always subgroups of $\mathbb{Z}$, they must then be trivial, and so $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular by Theorem A. ∎
Next we show that we can recover [20, Theorem 10], which states that the Leavitt path algebra of an arbitrary graph over a field is graded von Neumann regular. In fact, we can get away with having coefficients in a commutative regular ring. Every Leavitt path algebra $L_{K}(E)$ over a commutative ring $K$ is $\mathbb{Z}$-graded, with the grading given by homogeneous components
@@
Let $E$ be an arbitrary graph and $K$ a unital commutative von Neumann regular unital ring. Then the Leavitt path algebra $L_{K}(E)$ is a graded von Neumann regular ring.
@@
$\mathscr{G}_{0}(F,n)=\bigcup_{\alpha,\beta\in F^{*},\mathop{\boldsymbol{r}}\nolimits(\alpha)=\mathop{\boldsymbol{r}}\nolimits(\beta),|\alpha|=|\beta|\leq n}Z(\alpha,\emptyset,\beta).$